# Wasserstein Proximal Coordinate Gradient Algorithms

### v1

@@ 4.1 Convergence rates of WPCG under $\lambda$-QG assumption
Thus in our proof, different from the Euclidean case, we cannot directly calculate the norm of the blockwise Wasserstein gradient of the updated coordinate and take expectation with respect to the index of the updated coordinate due to the same non-tensorization issue described in (18). We instead prove a high probability bound of the norm of blockwise Wasserstein gradient to control the functional value of $\mathcal{F}$ after each iteration. When this bound does not hold, we directly use the non-increasing property of the functional value to bound $\mathcal{F}$. By tuning the number $M$ of updates in each iteration, we can derive the convergence rate of WPCG-R in Theorem 4.8. It is still an open question whether the iteration complexity in the theorem can be improved by applying other strategies. ∎
@@ 4.2 Convergence rates of WPCG without $\lambda$-QG assumption
Theorem 4.11 (Polynomial convergence rate of WPCG without $\lambda$-QG).
@@
Remark 4.12 (Comparison with existing results in the Euclidean case).
When the QG condition is not met for some $\lambda>0$, the functional value in a typical first-order optimization algorithm decreases with a rate no faster than $O(k^{-1})$ in the worst case. This has been demonstrated through various studies on the Euclidean space especially for the sequential update scheme. Theorem 3 in [50] shows that the function value decreases with a rate $O(k^{-1})$ when applying the coordinate GD with the sequential update scheme. Theorem 11.18 in [6] proves the same $O(k^{-1})$ convergence rate for the coordinate proximal GD with the sequential update scheme. Here, we derive the same convergence rate for WPCG with all three update schemes. ∎
The assumption of compactness of the parameter space is frequently utilized when applying the Wasserstein proximal gradient scheme to optimization problems, as seen in works by [53, 42]. This assumption plays a crucial role in our proof, allowing us to derive a uniform bound for $\operatorname{W}_{2}^{2}(\rho^{k},\rho^{\ast})$ and the first-order optimality condition of Wasserstein proximal gradient scheme (Lemma B.1). The compactness is specifically used to determine the first variation of $\operatorname{W}_{2}^{2}(\mu,\nu)$ with respect to $\mu$. More details on this can be found in Proposition 7.17 and Theorem 1.52 in [42]. It would be interesting to explore potential methods for relaxing this technical condition.
@@ 4.3 Convergence rates of inexact WPCG
$\eta_{j}^{k+1}=T_{\widetilde{\rho}_{j}^{k+1}}^{\widetilde{\rho}_{j}^{k}}-\textrm{Id}-\tau\bigg{[}\nabla V_{j}^{k}+\nabla h_{j}^{\prime}(\widetilde{\rho}_{j}^{k+1})+\nabla\!\int_{\mathcal{X}}W_{j}(\cdot,y)\,{\rm d}\widetilde{\rho}_{j}^{k+1}(y)+\nabla\!\int_{\mathcal{X}}W_{j}(y,\cdot)\,{\rm d}\widetilde{\rho}_{j}^{k+1}(y)\bigg{]}.$
as the first variation of the objective functional in the subproblem (13) evaluated at $\widetilde{\rho}_{j}^{k+1}$. If $\widetilde{\rho}_{j}^{k+1}$ is the exact solution of the subproblem (13), the first-order optimality condition (Lemma B.1) implies $\eta_{j}^{k+1}=0$. Therefore, we can employ $\|\eta_{j}^{k+1}\|_{L^{2}(\mathcal{X}_{j};\widetilde{\rho}_{j}^{k+1})}$ to characterize the numerical error incurred while solving the subproblem. Let $\{\varepsilon_{j}^{k}:j\in[m],k\in\mathbb{Z}_{+}\}$ be a sequence of error tolerance levels such that
@@
Theorem 4.13 (Convergence rate of inexact WPCG-P under $\lambda$-QG).
@@
Remark 4.14 (Impact of numerical error).
@@
Theorem 4.15 (Polynomial convergence rate of inexact WPCG-P without $\lambda$-QG).
@@ 5 Numerical experiments
In this section, we demonstrate the application of our WPCG algorithms to approximate the posterior distribution via the mean-field variational inference in Example 1.1 and to compute the stationary distribution of the multi-species systems with cross-interaction in Example 1.2.
@@ 5.1 Mean-field variational inference
where $\Pi_{n}$ is the posterior distribution with the density function given by (3). As a corollary of Theorem 4.2, we have the following convergence result about computing $\widehat{\pi}_{n}$ via WPCG-P.
Corollary 5.1.
@@
Note that in Corollary 5.1, both $\lambda_{n}$ and $L_{n}$ depend on the samples $X_{1},\dots,X_{n}$. When the log-likelihood function $\log p(x\,|\,\theta)$ satisfies certain mild conditions [e.g., Assumptions 2 and 3 in 34], the sample convexity parameter $\lambda_{n}$ and the smoothness parameter $L_{n}$ will concentration in the small neighborhoods of the population convexity parameter $\lambda=\mathbb{E}_{\theta^{\ast}}\lambda_{n}$ and the smoothness parameter $L=\mathbb{E}_{\theta^{\ast}}L_{n}$ respectively with high probability.
Remark 5.2.
When applying WPCG-P to MFVI, our method coincides with representations in [22]. However, Ghosh et al. [22] only studied the convergence of the piecewise constant interpolation $\{\rho_{t}=\rho^{\lfloor t/\tau\rfloor}:t\geq 0\}$ towards the solution of its corresponding WGF [Equation (15) in 22] as the step size $\tau\to 0$ when $V_{n}$ is strongly convex. Yao and Yang [53] proposed the MF-WGF algorithm for solving MFVI with latent variables numerically. They showed exponential convergence of iterates towards the mean-field approximation $\widehat{\pi}_{n}$ of the true posterior in $\operatorname{W}_{2}$ sense. Their proof heavily depended on the strong convexity of the negative log-likelihood function, while our analysis only uses the strong convexity to show QG condition. Lambert et al. [30] studied Gaussian variational inference (i.e., $m=1$ and the variational family is restricted to all Gaussian distributions) and showed that the corresponding gradient flow (so-called Bures–Wasserstein gradient flow) converges to the Gaussian variational approximation exponentially fast in $\operatorname{W}_{2}$ sense. Compared with Gaussian variational inference, our method is more flexible with the number of blocks and the variational family. ∎
@@ 5.2 Equilibrium in multi-species systems
Remark 5.4.
(i) In this lemma, we assume $K_{ij}=K_{ji}$ for all $i,j\in[m]$ so that the evolution of $\{\rho_{j}(\cdot,t):j\in[m],t\geq 0\}$ follows the WGF of some functional $\mathcal{F}$. When $m=2$, this condition can be relaxed to $K_{12}=\alpha K_{21}$ for some constant $\alpha>0$ by rescaling the WGF of $\rho_{1}$ and $\rho_{2}$ [18]. (ii) The choice of $h_{j}(x)=x\log x$ corresponds to the standard diffusion without medium of species $j$, and (5) turns to be the PDE of multi-species stochastic interacting particle systems [16]. As a comparison, $h_{j}(x)=x^{m_{j}}$ [^2] Here the exponent $m_{j}>1$ depends on physical properties of the diffusing material. corresponds to a porous medium type diffusion [36]. When $m=1$, (5) is the aggregation equation of interacting particle systems [11, 14, 32]; when $m=2$ and $V_{j}=h_{j}=0$, (5) turns into the PDE of non-local interaction systems with two species [18, 20]. To our knowledge, Equation (5) has not been studied for general $m\in\mathbb{N}^{\ast}$ and functions $h_{j}$, $j\in[m]$. We believe that our theory is also useful to study the existence and the uniqueness of this kind of PDEs from the perspective of WGF. ∎
In the numerical study, we consider the following test example [16, 26] with $m=3$ species in $\mathbb{R}^{2}$. We specify the functions
@@
Figure 4 shows the scatter plots of all species with different $\alpha$ and $\beta$. When $\alpha$ gets larger, all species concentrate more around the corresponding centers ($\theta_{1}$, $\theta_{2}$, and $\theta_{3}$) due to a larger external force. In contrast, larger $\beta$ makes all species less concentrate since the entropy term penalizes the concentration of species. Figure 5 shows the optimization error $\operatorname{W}_{2}^{2}(\rho_{\theta}^{k},\rho_{\theta}^{\ast})$ for different $\alpha$ and $\beta$ and the variance of the first variation when $\alpha=\beta=1$. In Figure 5a, the optimization error decays faster with larger $\alpha$ (compare the blue line with the orange line) since the convexity gets stronger, while changing $\beta$ will not change the convergence speed (compare the blue line with the green line) since the entropy term does not provide any convexity to the whole functional. Figure 5b plots the variance of the first variation $\frac{\delta\mathcal{F}}{\delta\rho_{j}}(\rho^{k})(X_{j}^{k})$ where $X_{j}^{k}\sim\rho_{j}^{k}$. This variance converges to 00 if and only if the first variation $\frac{\delta\mathcal{F}}{\delta\rho_{j}}(\rho^{k})$ converges to a constant [$\rho_{j}$]-a.e., which indicates that $(\rho_{1}^{k},\rho_{2}^{k},\rho_{3}^{k})$ converges to the minimum of $\mathcal{F}$.
Figure 6 shows the scatter plot of all species when there exists an extra super-quadratic term $\|x\|^{4}/4$ in the potential function $V$. Due to this super-quadratic term, all species get closer to the origin when the system is at its equilibrium. Figure 7 shows the optimization error $\operatorname{W}_{2}^{2}(\rho_{\theta}^{k},\rho_{\theta}^{\ast})$ and the variance of the first variation. Figure 7a compares the optimization error derived by WPCG with the FA approach and the SDE approach. Compare with the FA approach, the optimization error derived by applying the SDE approach (the blue line and the green line) is dominated by the approximation error after several iterations. Due to the existence of the super-quadratic term, a smaller step size has to be chosen to make the SDE approach converge when the initialization gets farther from the origin, which increases the number of iterations for the SDE approach to converge. In contrast, the optimization error derived by the FA approach (the red line and the orange line) is not affected by the super-quadratic term as predicted by our theoretical result. Fig 7b plots the variance of the first variation $\frac{\delta\mathcal{F}}{\delta\rho_{j}}(\rho^{k})$. The variance converges to zero as a sign of convergence of $\rho^{k}$ to the minimum of $\mathcal{F}$.
@@ 5.3 Real data example
We apply the WPCG-P algorithm to analyze the Pima Indian diabetes dataset [43]. This dataset comprises 8 medical predictors (pregnancies, glucose, blood pressure, skin thickness, insulin, body mass index, diabetes pedigree function, age) and a binary outcome variable (00 for no diabetes, $1$ have diabetes) from 768 individuals.
@@ 6 Conclusions and discussions
    Another potential topic is to consider the coordinate Wasserstein forward-backward gradient descent algorithm, which updates each coordinate by approximating the potential energy $\mathcal{V}$ in first order (forward step) followed by a proximal descent step (backward step). This algorithm corresponds to the proximal coordinate gradient descent algorithm in the Euclidean case. The case $m=1$ on the Wasserstein space has been studied by [41]. We believe our analysis of WPCG can be extended to the coordinate Wasserstein forward-backward gradient descent algorithm.
@@ Appendix A Proof of Main Results
A.1 Parallel update scheme: proof of Theorem 4.2
@@
Lemma A.1 (sufficient decrease).
@@
Lemma A.2 (bound of functional value).
@@
is defined in Theorem 4.2. Therefore, we have
@@
A.2 Sequential update scheme: proof of Theorem 4.5
@@
Lemma A.3 (sufficient decrease).
@@
A.3 Random update scheme: proof of Theorem 4.8
@@
Lemma A.5 (sufficient decrease).
@@
Lemma A.6 (bound of functional value).
@@
A.4 Case $\lambda=0$: proof of Theorem 4.11

### v2

@@ 4.1 Convergence Rates of WPCG under $\lambda$-QG Assumption
Thus in our proof, different from the Euclidean case, we cannot directly calculate the norm of the blockwise Wasserstein gradient of the updated coordinate and take expectation with respect to the index of the updated coordinate due to the same non-tensorization issue described in (18). We instead prove a high probability bound of the norm of blockwise Wasserstein gradient to control the functional value of $\mathcal{F}$ after each iteration. When this bound does not hold, we directly use the non-increasing property of the functional value to bound $\mathcal{F}$. By tuning the number $M$ of updates in each iteration, we can derive the convergence rate of WPCG-R in Theorem 15. It is still an open question whether the iteration complexity in the theorem can be improved by applying other strategies. ∎
@@ 4.2 Convergence Rates of WPCG without $\lambda$-QG Assumption
Theorem 18 (Polynomial convergence rate of WPCG without $\lambda$-QG).
@@
Remark 19 (Comparison with existing results in the Euclidean case).
When the QG condition is not met for some $\lambda>0$, the functional value in a typical first-order optimization algorithm decreases with a rate no faster than $O(k^{-1})$ in the worst case. This has been demonstrated through various studies on the Euclidean space especially for the sequential update scheme. Theorem 3 in (Wright, 2015) shows that the function value decreases with a rate $O(k^{-1})$ when applying the coordinate GD with the sequential update scheme. Theorem 11.18 in (Beck, 2017) proves the same $O(k^{-1})$ convergence rate for the coordinate proximal GD with the sequential update scheme. Here, we derive the same convergence rate for WPCG with all three update schemes. ∎
The assumption of compactness of the parameter space is frequently utilized when applying the Wasserstein proximal gradient scheme to optimization problems, as seen in works by Yao and Yang (2022); Santambrogio (2015). This assumption plays a crucial role in our proof, allowing us to derive a uniform bound for $\operatorname{W}_{2}^{2}(\rho^{k},\rho^{\ast})$ and the first-order optimality condition of Wasserstein proximal gradient scheme (Lemma 9). The compactness is specifically used to determine the first variation of $\operatorname{W}_{2}^{2}(\mu,\nu)$ with respect to $\mu$. More details on this can be found in Proposition 7.17 and Theorem 1.52 in (Santambrogio, 2015). It would be interesting to explore potential methods for relaxing this technical condition.
@@ 4.3 Convergence Rates of Inexact WPCG
$\eta_{j}^{k+1}\!=\!T_{\widetilde{\rho}_{j}^{k+1}}^{\widetilde{\rho}_{j}^{k}}\!-\textrm{Id}-\!\tau\bigg{[}\nabla V_{j}^{k}+\nabla h_{j}^{\prime}(\widetilde{\rho}_{j}^{k+1})+\nabla\!\int_{\mathcal{X}}W_{j}(\cdot,y)\,{\rm d}\widetilde{\rho}_{j}^{k+1}(y)+\nabla\!\int_{\mathcal{X}}W_{j}(y,\cdot)\,{\rm d}\widetilde{\rho}_{j}^{k+1}(y)\bigg{]}$
as the first variation of the objective functional in the subproblem (13) evaluated at $\widetilde{\rho}_{j}^{k+1}$. If $\widetilde{\rho}_{j}^{k+1}$ is the exact solution of the subproblem (13), the first-order optimality condition (Lemma 9) implies $\eta_{j}^{k+1}=0$. Therefore, we can employ $\|\eta_{j}^{k+1}\|_{L^{2}(\mathcal{X}_{j};\widetilde{\rho}_{j}^{k+1})}$ to characterize the numerical error incurred while solving the subproblem. Let $\{\varepsilon_{j}^{k}:j\in[m],k\in\mathbb{Z}_{+}\}$ be a sequence of error tolerance levels such that
@@
Theorem 20 (Convergence rate of inexact WPCG-P under $\lambda$-QG).
@@
Remark 21 (Impact of numerical error).
@@
Theorem 22 (Polynomial convergence rate of inexact WPCG-P without $\lambda$-QG).
@@ 5 Numerical Experiments
In this section, we demonstrate the application of our WPCG algorithms to approximate the posterior distribution via the mean-field variational inference in Example 1 and to compute the stationary distribution of the multi-species systems with cross-interaction in Example 2.
@@ 5.1 Mean-field Variational Inference
where $\Pi_{n}$ is the posterior distribution with the density function given by (3). As a corollary of Theorem 9, we have the following convergence result about computing $\widehat{\pi}_{n}$ via WPCG-P.
Corollary 23.
@@
Note that in Corollary 23, both $\lambda_{n}$ and $L_{n}$ depend on the samples $X_{1},\dots,X_{n}$. When the log-likelihood function $\log p(x\,|\,\theta)$ satisfies certain mild conditions (e.g., Assumptions 2 and 3 in Mei et al., 2016), the sample convexity parameter $\lambda_{n}$ and the smoothness parameter $L_{n}$ will concentration in the small neighborhoods of the population convexity parameter $\lambda=\mathbb{E}_{\theta^{\ast}}\lambda_{n}$ and the smoothness parameter $L=\mathbb{E}_{\theta^{\ast}}L_{n}$ respectively with high probability.
Remark 24.
When applying WPCG-P to MFVI, our method coincides with representations in (Ghosh et al., 2022). However, Ghosh et al. (2022) only studied the convergence of the piecewise constant interpolation $\{\rho_{t}=\rho^{\lfloor t/\tau\rfloor}:t\geq 0\}$ towards the solution of its corresponding WGF (Equation (15) in Ghosh et al., 2022) as the step size $\tau\to 0$ when $V_{n}$ is strongly convex. Yao and Yang (2022) proposed the MF-WGF algorithm for solving MFVI with latent variables numerically. They showed exponential convergence of iterates towards the mean-field approximation $\widehat{\pi}_{n}$ of the true posterior in $\operatorname{W}_{2}$ sense. Their proof heavily depended on the strong convexity of the negative log-likelihood function, while our analysis only uses the strong convexity to show QG condition. Lambert et al. (2022) studied Gaussian variational inference (i.e., $m=1$ and the variational family is restricted to all Gaussian distributions) and showed that the corresponding gradient flow (so-called Bures–Wasserstein gradient flow) converges to the Gaussian variational approximation exponentially fast in $\operatorname{W}_{2}$ sense. Compared with Gaussian variational inference, our method is more flexible with the number of blocks and the variational family. ∎
@@ 5.2 Equilibrium in Multi-species Systems
Remark 26.
(i) In this lemma, we assume $K_{ij}=K_{ji}$ for all $i,j\in[m]$ so that the evolution of $\{\rho_{j}(\cdot,t):j\in[m],t\geq 0\}$ follows the WGF of some functional $\mathcal{F}$. When $m=2$, this condition can be relaxed to $K_{12}=\alpha K_{21}$ for some constant $\alpha>0$ by rescaling the WGF of $\rho_{1}$ and $\rho_{2}$ (Di Francesco and Fagioli, 2013). (ii) The choice of $h_{j}(x)=x\log x$ corresponds to the standard diffusion without medium of species $j$, and (5) turns to be the PDE of multi-species stochastic interacting particle systems (Daus et al., 2022). As a comparison, $h_{j}(x)=x^{m_{j}}$ [^2] Here the exponent $m_{j}>1$ depends on physical properties of the diffusing material. corresponds to a porous medium type diffusion (Otto, 2001). When $m=1$, (5) is the aggregation equation of interacting particle systems (Cabrales et al., 2020; Carrillo et al., 2019; Li and Rodrigo, 2010); when $m=2$ and $V_{j}=h_{j}=0$, (5) turns into the PDE of non-local interaction systems with two species (Di Francesco and Fagioli, 2013; Evers et al., 2017). To our knowledge, Equation (5) has not been studied for general $m\in\mathbb{N}^{\ast}$ and functions $h_{j}$, $j\in[m]$. We believe that our theory is also useful to study the existence and the uniqueness of this kind of PDEs from the perspective of WGF. ∎
In the numerical study, we consider the following test example (Daus et al., 2022; Jin et al., 2020) with $m=3$ species in $\mathbb{R}^{2}$. We specify the functions
@@
Figure 4 shows the scatter plots of all species with different $\alpha$ and $\beta$. When $\alpha$ gets larger, all species concentrate more around the corresponding centers ($\theta_{1}$, $\theta_{2}$, and $\theta_{3}$) due to a larger external force. In contrast, larger $\beta$ makes all species less concentrate since the entropy term penalizes the concentration of species. Figure 5 shows the optimization error $\operatorname{W}_{2}^{2}(\rho_{\theta}^{k},\rho_{\theta}^{\ast})$ for different $\alpha$ and $\beta$ and the variance of the first variation when $\alpha=\beta=1$. In Figure 5(a), the optimization error decays faster with larger $\alpha$ (compare the blue line with the orange line) since the convexity gets stronger, while changing $\beta$ will not change the convergence speed (compare the blue line with the green line) since the entropy term does not provide any convexity to the whole functional. Figure 5(b) plots the variance of the first variation $\frac{\delta\mathcal{F}}{\delta\rho_{j}}(\rho^{k})(X_{j}^{k})$ where $X_{j}^{k}\sim\rho_{j}^{k}$. This variance converges to 00 if and only if the first variation $\frac{\delta\mathcal{F}}{\delta\rho_{j}}(\rho^{k})$ converges to a constant [$\rho_{j}$]-a.e., which indicates that $(\rho_{1}^{k},\rho_{2}^{k},\rho_{3}^{k})$ converges to the minimum of $\mathcal{F}$.
Figure 6 shows the scatter plot of all species when there exists an extra super-quadratic term $\|x\|^{4}/4$ in the potential function $V$. Due to this super-quadratic term, all species get closer to the origin when the system is at its equilibrium. Figure 7 shows the optimization error $\operatorname{W}_{2}^{2}(\rho_{\theta}^{k},\rho_{\theta}^{\ast})$ and the variance of the first variation. Figure 7(a) compares the optimization error derived by WPCG with the FA approach and the SDE approach. Compare with the FA approach, the optimization error derived by applying the SDE approach (the blue line and the green line) is dominated by the approximation error after several iterations. Due to the existence of the super-quadratic term, a smaller step size has to be chosen to make the SDE approach converge when the initialization gets farther from the origin, which increases the number of iterations for the SDE approach to converge. In contrast, the optimization error derived by the FA approach (the red line and the orange line) is not affected by the super-quadratic term as predicted by our theoretical result. Fig 7(b) plots the variance of the first variation $\frac{\delta\mathcal{F}}{\delta\rho_{j}}(\rho^{k})$. The variance converges to zero as a sign of convergence of $\rho^{k}$ to the minimum of $\mathcal{F}$.
@@ 5.3 Real Data Example
We apply the WPCG-P algorithm to analyze the Pima Indian diabetes data set (Smith et al., 1988). This data set comprises 8 medical predictors (pregnancies, glucose, blood pressure, skin thickness, insulin, body mass index, diabetes pedigree function, age) and a binary outcome variable (00 for no diabetes, $1$ have diabetes) from 768 individuals.
@@ 6 Conclusions and Discussions
    Another potential topic is to consider the coordinate Wasserstein forward-backward gradient descent algorithm, which updates each coordinate by approximating the potential energy $\mathcal{V}$ in first order (forward step) followed by a proximal descent step (backward step). This algorithm corresponds to the proximal coordinate gradient descent algorithm in the Euclidean case. The case $m=1$ on the Wasserstein space has been studied by Salim et al. (2020). We believe our analysis of WPCG can be extended to the coordinate Wasserstein forward-backward gradient descent algorithm.
@@
Acknowledgments and Disclosure of Funding
Xiaohui Chen was partially supported by NSF CAREER grant DMS-2347760, NSF grant DMS-2413404, and a gift from the Simons Foundation. Yun Yang was partially supported by NSF grant DMS-2210717.
@@ Appendix A Proof of Main Results
A.1 Parallel Update Scheme: Proof of Theorem 9
@@
Lemma 1 (sufficient decrease).
@@
Lemma 2 (bound of functional value).
@@
is defined in Theorem 9. Therefore, we have
@@
A.2 Sequential Update Scheme: Proof of Theorem 12
@@
Lemma 3 (sufficient decrease).
@@
A.3 Random Update Scheme: Proof of Theorem 15
@@
Lemma 5 (sufficient decrease).
@@
Lemma 6 (bound of functional value).
@@
A.4 Case $\lambda=0$: Proof of Theorem 18